# Tight Guarantees for Cut-Relative Survivable Network Design via a Decomposition Technique

### v2

@@ 4 Structure of feasible $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ solutions: a decomposition result
In the splitting operation, we pick a small cut $A$ of maximum deficiency, and move to $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ instances on the smaller graphs $G[{A}]$, and $G[{\overline{A}}]$ obtained by restricting $f$ to $A$, and restricting $f$ to $\overline{A}$ respectively (Theorem 4.7). Lemma 4.5 states some basic properties of the restriction of $f$. Part (b) will be quite useful in proving that the splitting operation maintains feasibility, and part (a) allows us to repeat the splitting operation on the smaller instances. For a set $S\subseteq V$, and set $F\subseteq E$ of edges, we use $F(S)$ to denote the edges of $F$ with both ends in $S$.
@@
For part (a), we can proceed in two ways. As noted earlier, $f_{S}$ is the symmetric, normalized version $f^{\prime}$, so part (a) follows from Lemma 2.1 (a). Alternatively, observe that $f_{S}$ is the projection of $f^{\prime}$ onto $S$ with respect to $\mathcal{C}:=\{\emptyset,\overline{S}\}$, which is closed under set intersections, unions, differences, and complements with respect to $V-S$. So by Lemma 2.1 (c), we obtain that $f_{S}$ is weakly supermodular and symmetric. Also, $f_{S}$ is normalized by definition.
@@
$\begin{split}\operatorname{def}_{f,F}(T)+\operatorname{def}_{f,F}(A)&\leq\operatorname{def}_{f,F}(T\cap A)+\operatorname{def}_{f,F}(T\cup A)=\operatorname{def}_{f,F}(T\cap A)+\operatorname{def}_{f,F}(\overline{A}-T)\ \ \text{or}\\ \operatorname{def}_{f,F}(T)+\operatorname{def}_{f,F}(A)&\leq\operatorname{def}_{f,F}(A-T)+\operatorname{def}_{f,F}(T-A)\end{split}\vspace*{-2ex}$
@@ 5 Algorithmic implication: LP-relative $2$-approximation algorithm
[…]
@@
[…]
@@
Throughout this section, $f$ and $G$ always denote the base cut-requirement function $f$ and graph $G$ that are given as input to $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$. Theorem 4.1 does not pinpoint a specific decomposition, and only mentions some properties that the $(f_{i},V_{i})$ tuples should satisfy. By “the decomposition in Theorem 4.1,” we mean the decomposition given by its proof, wherein we obtain a maximum deficiency set $A\subseteq V$ as in Theorem 4.7, and recurse on the $\mathsf{CR}\text{-}{(f_{A},G[{A}])}\text{-}\mathsf{NDP}$ and $\mathsf{CR}\text{-}{(f_{\overline{A}},G[{\overline{A}}])}\text{-}\mathsf{NDP}$-instances. This process can be conveniently described in terms of a decomposition tree as follows. Each node of the tree is labeled by a tuple $(h,S)$, where $S\subseteq V$ and $g:2^{S}\mapsto\mathbb{Z}$ is a weakly-supermodular, symmetric, normalized cut-requirement function on $S$; we will often simply say node $(h,S)$. The root of the tree is labeled $(f,V)$. For each (current) leaf node $(h,S)$, we find $A\subseteq S$ such that $\operatorname{def}_{h,G[{S}]}(A)=\max_{T\subseteq S}\operatorname{def}_{h,G[{S}]}(T)$. If $\operatorname{def}_{f,G[{S}]}(A)\leq 0$, then there is no small-$(h,G[{S}])$-cut and $(h,S)$ gets fixed as a leaf node. Otherwise, we create two children of $(h,S)$ labeled $(h_{1},A)$ and $(h_{2},S-A)$, where $h_{1}$ and $h_{2}$ are the restrictions of $h$ to $A$ and $S-A$ respectively (see Definition 4.4), and continue. We call this tree the canonical decomposition tree, and refer to the decomposition given by its leaves as the canonical decomposition.
@@
Another consequence of the properties we establish is that they allow us to succinctly describe the $(f_{i},V_{i})$ tuples given by the leaves of the canonical decomposition tree (see Lemmas 6.6 and 6.8). While this yields a self-contained, succinct description of the canonical decomposition, we emphasize that in order to prove that the ${f}\text{-}\mathsf{NDP}$-instances given by this succinct description satisfy the properties stated in Theorem 4.1, we need to go via the canonical decomposition and exploit Lemma 6.1 (and thus our approach of iteratively simplifying the instance via splitting along maximum-deficiency small cuts). So having this succinct description does not simplify the underlying arguments; see also Section 6.1.1.
@@
    $\max\,\bigl{\{}\operatorname{def}_{h,G[{S}]}(T):\ T\subseteq S\text{ is an $s$-$t$ cut}\bigr{\}}=\max\,\bigl{\{}\operatorname{def}_{f,G}(T):\ T\subseteq V\text{ is an $s$-$t$ cut}\bigr{\}}.$
@@
We now prove (8). We have $\operatorname{def}_{h_{i},G[{S_{i}}]}(X_{i})+\operatorname{def}_{h_{i},G[{S_{i}}]}(S_{i-1})\leq\operatorname{def}_{h_{i},G[{S_{i}}]}(U_{i-1})+\operatorname{def}_{h_{i},G[{S_{i}}]}(Y_{i-1})$ where $U_{i-1}=X_{i}\cap S_{i-1}$, $Y_{i-1}=X_{i}\cup S_{i-1}$, or $U_{i-1}=S_{i-1}-X_{i}$, $Y_{i-1}=X_{i}-S_{i-1}$. By the definition of $S_{i-1}$, we have $\operatorname{def}_{h_{i},G[{S_{i}}]}(S_{i-1})\geq\operatorname{def}_{h_{i},G[{S_{i}}]}(Y)$ for all $Y\subseteq S_{i}$. So the above inequality yields that $\operatorname{def}_{h_{i},G[{S_{i}}]}(X_{i})\leq\operatorname{def}_{h_{i},G[{S_{i}}]}(U_{i-1})$. Also, by part (b), we have $\operatorname{def}_{h_{i},G[{S_{i}}]}(U_{i-1})\leq\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(U_{i-1})$. We know that $U_{i-1}$ is $X_{i}\cap S_{i-1}=X_{i-1}$ or $S_{i-1}-X_{i}$. So since $\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(\cdot)$ is symmetric, we therefore obtain that $\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(U_{i-1})=\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(X_{i}\cap S_{i-1})=\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(X_{i-1})$. Combining everything, we obtain that (8) holds for index $i$.
@@
Furthermore, note that if can solve weighted max $s$-$t$ deficient cut in polytime, then for any leaf $(f_{i},S)$ of the canonical decomposition tree, we can give an efficient separation oracle for the LP $(\text{{NDP}}_{f_{i},G[{S}]})$. This is because exactly as in Lemma 6.1 (c), we have $\max\,\bigl{\{}\operatorname{wdef}_{f_{i},G[{S}]}(T):\ T\subseteq S\text{ is an $s$-$t$ cut}\bigr{\}}=\max\,\bigl{\{}\operatorname{wdef}_{f,G}(T):\ T\subseteq V\text{ is an $s$-$t$ cut}\bigr{\}}$, and if $T$ attains the maximum on the right, then $T\cap S$ attains the maximum on the left. So we can find $\max_{T\subseteq S}\bigl{(}f_{i}(T)-x\bigl{(}\delta_{G[{S}]}(T)\bigr{)}$ by performing $|S|$ weighted max-$s$-$t$-deficient-cut computations (with respect to $(f,G)$).
@@ 6.1 Succinct description of the canonical decomposition
As noted earlier, one structural consequence of Lemma 6.1 is that one can give a succinct description of the $(f_{i},V_{i})$ tuples obtained at the leaves of the canonical decomposition tree. We show that the $V_{1},\ldots,V_{k}$ leaf-set partition is precisely the partition given by the (node-sets of the) components of $G-Z$, where $Z:=\bigcup_{S\subseteq V:\operatorname{def}_{f,G}(S)>0}\delta_{G}(S)$ is the set of all edges on the boundary of some small cut. We can also succinctly specify the $f_{i}:2^{V_{i}}\mapsto\mathbb{Z}$ functions for $i\in[k]$ (Lemma 6.8).
@@
The leaf sets $V_{1},\ldots,V_{k}$ of the canonical decomposition correspond to the node-sets of the components of $G-Z$.
@@
We first argue that the components of $G-Z$ yield a refinement of the $V_{1},\ldots,V_{k}$ partition. To do so, consider any node $(h,S)\neq(f,V)$ of the canonical decomposition tree, and let $(h_{1},S_{1})$ be the parent of $(h,S)$. We claim that $\delta_{G[{S_{1}}]}(S)\subseteq Z$. Applying this to every node of the canonical decomposition tree yields that $\delta_{G}(V_{i})\subseteq Z$ for every leaf-set $V_{i}$. To see the claim, applying part (a) of Lemma 6.1 to $(h_{1},S_{1})$ taking $T=S$, since $\operatorname{def}_{h_{1},G[{S_{1}}]}(S)>0$, we obtain that there is some small-$(f,G)$-cut $A\subseteq V$ such that $S=A\cap S_{1}$. Thus, $\delta_{G[{S_{1}}]}(S)\subseteq\delta_{G}(A)\subseteq Z$.
We next argue that no edge of $Z$ is internal to a leaf set, which will finish the proof. Suppose not, and there is some $e=uv\in Z$ with both $u,v\in V_{i}$ for some $i\in[k]$. Then, by the definition of $Z$, there is some small-$(f,G)$-cut $T$ such $e\in\delta_{G}(T)$. But then taking $s=u,t=v$, and by Lemma 6.1 (c), there is some small-$(f_{i},G[{V_{i}}])$-cut $T^{\prime}\subsetneq V_{i}$ separating $s$ and $t$. This yields a contradiction since $(f_{i},V_{i})$ is a leaf of the canonical decomposition tree.
We point out that for a leaf-set $V_{i}$, while $\delta_{G}(V_{i})\subseteq Z$, it need not be that $V_{i}$ is a small-$(f,G)$-cut. For instance, if the base cut-requirement function $f$ models an $s$-$t$ connectivity problem, then every small-$(f,G)$-cut must be an $s$-$t$ cut, and so unless there are only two leaf sets, this will not hold.
@@
6.1.1 Some remarks and discussion

### v3

@@ 4 Structure of feasible $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ solutions: a decomposition result
In the splitting operation, we pick a small cut $A$ of maximum deficiency, and move to $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$ instances on the smaller graphs G[A]G[{A}], and $G[{\overline{A}}]$ obtained by restricting $f$ to $A$, and restricting $f$ to $\overline{A}$ respectively (Theorem 4.7). Lemma 4.5 states some basic properties of the restriction of $f$. Part (b) will be quite useful in proving that the splitting operation maintains feasibility, and part (a) allows us to repeat the splitting operation on the smaller instances. For a set $S\subseteq V$, and set $F\subseteq E$ of edges, we use $F(S)$ to denote the edges of $F$ with both ends in $S$.
@@
For part (a), we can proceed in two ways. As noted earlier, $f_{S}$ is the symmetric, normalized version $f^{\prime}$, so part (a) follows from Lemma 2.1 (a). Alternatively, observe that $f_{S}$ is the projection of $f^{\prime}$ onto $S$ with respect to $\mathcal{C}:=\{\emptyset,\overline{S}\}$, which is closed under set intersections, unions, differences, and complements with respect to VSV-S. So by Lemma 2.1 (c), we obtain that $f_{S}$ is weakly supermodular and symmetric. Also, $f_{S}$ is normalized by definition.
@@
$\begin{split}\operatorname{def}_{f,F}(T)+\operatorname{def}_{f,F}(A)&\leq\operatorname{def}_{f,F}(T\cap A)+\operatorname{def}_{f,F}(T\cup A)=\operatorname{def}_{f,F}(T\cap A)+\operatorname{def}_{f,F}(\overline{A}-T)\ \ \text{or}\\ \operatorname{def}_{f,F}(T)+\operatorname{def}_{f,F}(A)&\leq\operatorname{def}_{f,F}(A-T)+\operatorname{def}_{f,F}(T-A)\end{split}\vskip-8.61108pt$
@@
Let $\mathit{OPT}$ denote the optimal value of $(\text{{CRLP}}_{f,G})$. Algorithm CRNDP-Alg returns a feasible $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$-solution $F\subseteq E$ such that $c(F)\leq 2\cdot\mathit{OPT}$.
@@
By Corollary 4.2, at least one edge gets added to $F$ in every iteration, so the algorithm terminates in at most $|E|$ iterations. The cost bound follows by induction on the number of iterations to termination. The base case is trivial. Suppose $F=Z_{1}\cup F^{\prime}$, where $F^{\prime}$ is the solution found in subsequent iterations for the instance $\mathsf{CR}\text{-}{(f^{\prime},G^{\prime})}\text{-}\mathsf{NDP}$. Let $\mathit{OPT}^{\prime}$ be the optimal solution to $(\text{{CRLP}}_{G^{\prime},f^{\prime}})$. Then, by induction, we have $c(F_{1})\leq 2\cdot\mathit{OPT}^{\prime}$ and $\mathit{OPT}^{\prime}\leq\sum_{e\in E^{\prime}}c_{e}\widehat{x}_{e}$, since $(\widehat{x}_{e})_{e\in E^{\prime}}$ is a feasible solution to $(\text{{CRLP}}_{f^{\prime},G^{\prime}})$. We also have $c(Z_{1})\leq 2\cdot\sum_{e\in Z_{1}}c_{e}\widehat{x}_{e}$. Therefore, $c(F)\leq 2c^{T}\widehat{x}=2\cdot\mathit{OPT}$.
@@
Throughout this section, $f$ and $G$ always denote the base cut-requirement function $f$ and graph $G$ that are given as input to $\mathsf{CR}\text{-}{f}\text{-}\mathsf{NDP}$. Theorem 4.1 does not pinpoint a specific decomposition, and only mentions some properties that the $(f_{i},V_{i})$ tuples should satisfy. By “the decomposition in Theorem 4.1,” we mean the decomposition given by its proof, wherein we obtain a maximum deficiency set $A\subseteq V$ as in Theorem 4.7, and recurse on the $\mathsf{CR}\text{-}{(f_{A},G[{A}])}\text{-}\mathsf{NDP}$ and $\mathsf{CR}\text{-}{(f_{\overline{A}},G[{\overline{A}}])}\text{-}\mathsf{NDP}$-instances. This process can be conveniently described in terms of a decomposition tree as follows. Each node of the tree is labeled by a tuple $(h,S)$, where $S\subseteq V$ and $h:2^{S}\mapsto\mathbb{Z}$ is a weakly-supermodular, symmetric, normalized cut-requirement function on $S$; we will often simply say node $(h,S)$. The root of the tree is labeled $(f,V)$. For each (current) leaf node $(h,S)$, we find $A\subseteq S$ such that $\operatorname{def}_{h,G[{S}]}(A)=\max_{T\subseteq S}\operatorname{def}_{h,G[{S}]}(T)$. If $\operatorname{def}_{f,G[{S}]}(A)\leq 0$, then there is no small-(h,G[S])(h,G[{S}])-cut and $(h,S)$ gets fixed as a leaf node. Otherwise, we create two children of $(h,S)$ labeled $(h_{1},A)$ and $(h_{2},S-A)$, where $h_{1}$ and $h_{2}$ are the restrictions of $h$ to $A$ and SAS-A respectively (see Definition 4.4), and continue. We call this tree the canonical decomposition tree, and refer to the decomposition given by its leaves as the canonical decomposition.
@@
Another consequence of the properties we establish is that they allow us to succinctly describe the $(f_{i},V_{i})$ tuples given by the leaves of the canonical decomposition tree (see Lemmas 6.6 and 6.8), as also a closely-related decomposition satisfying the properties stated in Theorem 4.1. While this yields a self-contained, succinct description of the canonical decomposition, we emphasize that in order to prove that the ${f}\text{-}\mathsf{NDP}$-instances given by this succinct description satisfy the properties stated in Theorem 4.1, we need to go via the canonical decomposition and exploit Lemma 6.1 (and thus our approach of iteratively simplifying the instance via splitting along maximum-deficiency small cuts). So having this succinct description does not simplify the underlying arguments; see also Section 6.1.1.
@@
[…]
@@
We now prove (8). We have $\operatorname{def}_{h_{i},G[{S_{i}}]}(X_{i})+\operatorname{def}_{h_{i},G[{S_{i}}]}(S_{i-1})\leq\operatorname{def}_{h_{i},G[{S_{i}}]}(U_{i-1})+\operatorname{def}_{h_{i},G[{S_{i}}]}(Y_{i-1})$ where $U_{i-1}=X_{i}\cap S_{i-1}$, $Y_{i-1}=X_{i}\cup S_{i-1}$, or $U_{i-1}=S_{i-1}-X_{i}$, $Y_{i-1}=X_{i}-S_{i-1}$. By the definition of $S_{i-1}$, we have $\operatorname{def}_{h_{i},G[{S_{i}}]}(S_{i-1})\geq\operatorname{def}_{h_{i},G[{S_{i}}]}(Y)$ for all $Y\subseteq S_{i}$. So the above inequality yields that $\operatorname{def}_{h_{i},G[{S_{i}}]}(X_{i})\leq\operatorname{def}_{h_{i},G[{S_{i}}]}(U_{i-1})$. Also, by part (b), we have $\operatorname{def}_{h_{i},G[{S_{i}}]}(U_{i-1})\leq\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(U_{i-1})$. We know that $U_{i-1}$ is $X_{i}\cap S_{i-1}=X_{i-1}$ or $S_{i-1}-X_{i}$. So since $\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(\cdot)$ is symmetric, we therefore obtain that $\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(U_{i-1})=\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(X_{i}\cap S_{i-1})=\operatorname{def}_{h_{i-1},G[{S_{i-1}}]}(X_{i-1})$. Combining everything, we obtain that (8) holds for index ii.
@@
[…]
@@ 6.1 Succinct description of the canonical decomposition
As noted earlier, one structural consequence of Lemma 6.1 is that one can give a succinct description of the $(f_{i},V_{i})$ tuples obtained at the leaves of the canonical decomposition tree. This also leads to an alternate decomposition satisfying the properties stated in Theorem 4.1. Define $Z:=\bigcup_{S\subseteq V:\operatorname{def}_{f,G}(S)>0}\delta_{G}(S)$ to be the set of all edges on the boundary of some small-$(f,G)$-cut; we sometimes refer to $Z$ as small-cut edges. Let $C_{1},\ldots,C_{p}$ be the node-sets of the components of GZG-Z. Note that for a component $C_{i}$, while $\delta_{G}(C_{i})\subseteq Z$, it need not be that $C_{i}$ is a small-$(f,G)$-cut. (For instance, if the base cut-requirement function $f$ models an $s$-$t$ connectivity problem, then every small-$(f,G)$-cut must be an $s$-$t$ cut; so unless there are only two small cuts, it will not be that every $C_{i}$ set is a small cut.)
Say that components $C_{i}$ and $C_{j}$ are non-separable if for every small-$(f,G)$-cut AA, we have that $C_{i}\subseteq A$ iff $C_{j}\subseteq A$. Clearly, non-separability is an equivalence relation. We show that the $V_{1},\ldots,V_{k}$ partition is precisely the partition formed by the equivalence classes of the non-separability relation. We can also succinctly specify the $f_{i}:2^{V_{i}}\mapsto\mathbb{Z}$ functions for $i\in[k]$ (Lemma 6.8).
@@
The leaf sets $V_{1},\ldots,V_{k}$ of the canonical decomposition are the equivalence classes of the non-separability relation. That is, for each $V_{i}$, we have a set $\mathsf{EQ}_{i}\subseteq[p]$ of indices such that: (i) $V_{i}=\bigcup_{j\in\mathsf{EQ}_{i}}C_{j}$; and (ii) for any two indices $j\in\mathsf{EQ}_{i}$, $j^{\prime}\in[p]$, components $C_{j}$ and $C_{j^{\prime}}$ are non-separable iff $j^{\prime}\in\mathsf{EQ}_{i}$. In particular, note that there are no edges of $Z$ internal to a leaf-set.
@@
We first argue that the $C_{1},\ldots,C_{p}$ partition is a refinement of the $V_{1},\ldots,V_{k}$ partition. To do so, we need to show that $\delta_{G}(V_{i})\subseteq Z$ for every leaf-set $V_{i}$. Consider any node $(h,S)\neq(f,V)$ of the canonical decomposition tree, and let $(h_{1},S_{1})$ be the parent of $(h,S)$. Applying part (a) of Lemma 6.1 to $(h_{1},S_{1})$ taking $T=S$, since $\operatorname{def}_{h_{1},G[{S_{1}}]}(S)>0$, we obtain that there is some small-$(f,G)$-cut $A\subseteq V$ such that $S=A\cap S_{1}$. Thus, $\delta_{G[{S_{1}}]}(S)\subseteq\delta_{G}(A)\subseteq Z$. Now for any leaf-set $V_{i}$, every edge in $\delta_{G}(V_{i})$ belongs to $\delta_{G[{S_{1}}]}(S)$, where $(h,S)$ is some ancestor of $(f_{i},V_{i})$ and $(h_{1},S_{1})$ is its parent. So we obtain that $\delta_{G}(V_{i})\subseteq Z$ for every leaf-set $V_{i}$.
Now, consider some leaf-set $V_{i}$, and consider any two components $C_{j},C_{j^{\prime}}$, where $C_{j}\subseteq V_{i}$. We have to show that $C_{j}$ and $C_{j^{\prime}}$ are non-separable iff $C_{j^{\prime}}\subseteq V_{i}$. Suppose first that $C_{j^{\prime}}\nsubseteq V_{i}$. Let $C_{j^{\prime}}\subseteq V_{q}$ for some leaf-set $V_{q}\neq V_{i}$. Let $(h,S)$ be the least common ancestor of $(f_{i},V_{i})$ and $(f_{q},V_{q})$, and let $(h_{1},A)$, $(h_{2},S-A)$ be the children of $(h,S)$. Clearly, $A$ separates $V_{i}$ and $V_{q}$. We have that $0<\operatorname{def}_{h,G[{S}]}(A)=\max\,\{\operatorname{def}_{h,G[{S}]}(T):T\subseteq S\}$. So for any $s\in A$, $t\in S-A$, by Lemma 6.1 (c), we may assume that $A=T\cap S$, where $\operatorname{def}_{f,G}(T)=\operatorname{def}_{h,G[{S}]}(A)$. It follows that $T$ is a small-$(f,G)$-cut that separates $V_{i}$ and $V_{q}$, and hence also separates $C_{j}$ and $C_{j^{\prime}}$.
Now suppose that $C_{j^{\prime}}\subseteq V_{i}$. Suppose there is some small-$(f,G)$-cut $T$ that separates $C_{j}$ and $C_{j^{\prime}}$. Then taking any node $s\in C_{j}$ and any $t\in C_{j^{\prime}}$, by Lemma 6.1 (c), there is some small-$(f_{i},G[{V_{i}}])$-cut $T^{\prime}\subsetneq V_{i}$ separating $s$ and $t$. This yields a contradiction since $(f_{i},V_{i})$ is a leaf of the canonical decomposition tree.
@@
6.1.1 Further remarks and discussion
Alternate “direct” proof.